Alfredo is 5 years older than john and Maria is 4 years younger than john. If the of their ages is 34,what equation can be used to find John's age?
Found 2 solutions by ankor@dixie-net.com, Edwin McCravy:
Answer by ankor@dixie-net.com(22740)

Alfredo is 5 years older than john
a = j + 5
and Maria is 4 years younger than john.
m = j - 4
If the sum of their ages is 34,
a + j + m = 34
what equation can be used to find John's age?
replace a with (j+5), replace m with (j-4)
(j+5) + j + (j-4) = 34
combine like terms and find j

Answer by Edwin McCravy(20060)

Alfredo is 5 years older than john
John's age = J
Alfredo's age = J+5

and Maria is 4 years younger than john.
Maria's age = J-4

If the SUM of their ages is 34,
John's age + Alfredo's age + Maria's age = 34
     J     +      J+5      +      J-4    = 34

what equation can be used to find John's age?
Get it all together:

J + J+5 + J-4 = 34
         3J+1 = 34

You finish.

Edwin